OpenStudy (anonymous):

2(4r+4)=-16 please help

OpenStudy (kirbykirby):

8r+8=-16 8r=-24 r=-24/8 r=-3

OpenStudy (anonymous):

thankyou:)

OpenStudy (kirbykirby):

:)